# Context-aware knowledge graph framework for traffic speed forecasting using graph neural network

### v1

@@ IV-E Dual-view feature importance
Figure 3(a) demonstrates the heatmap derived from context-view attention weights in Equation 12 to investigate the contribution of context features. For the spatial unit, the prominence of the relation-dependent road feature at $[1]_{a}$ stands out, surpassing the influence of the road-only spatial feature at $[0]_{a}$ and the road-only temporal feature at $[10]_{a}$. This phenomenon suggests that isolated road representations without additional context information are less predictive. Moreover, the weights in Rows $[4-8]_{a}$ are notably elevated, indicating that spatial links—ranging from proximate to extended connections—play a crucial role in capturing the spatial dependency affecting traffic prediction. These links also exhibit a strong interplay with temporal features, particularly time indicators at Col $[11]_{a}$ and weather conditions at Col $[13]_{a}$.
@@
Regarding the temporal unit, all features within $[11-13]_{a}$ manifest as critical. First, time indicators and weather conditions in Rows $[11,13]_{a}$ demonstrate substantial weights individually and mutually. Their interaction with the relation-dependent road feature at Col $[1]_{a}$ is also of great importance. Second, the jam factor at $[12]_{a}$ stands out for its significant self-weight, aligning with the intuitive understanding that the congestion level is directly predictive of traffic speed with an immediate impact. Furthermore, hourly and weekly patterns in Rows $[14,16]_{a}$ carry high weights, whereas daily patterns in Row $[15]_{a}$ appear less influential. The prominence of hourly patterns can be attributed to their immediate relevance to the prediction timeframe. Meanwhile, weekly patterns demonstrate consistent regularity, capturing the cyclical nature of traffic flow where behaviors on a specific day, like Monday, tend to mirror those from the previous week. In contrast, daily patterns exhibit a weaker correlation, suggesting that the connection between consecutive days, such as Friday to the following Monday, possesses limited predictive value in this context.
@@ V-A CKG embedding analysis
Regarding $\mathbb{K}_{S}$, we have several findings. (1) Head and tail predictions. An imbalance in MR values for left-side and right-side evaluations is observed for both methods in Figure 4(a, b), with the both-side case presenting intermediate values. This imbalance reveals that predicting the head entity (e.g., $poiType$ and $landType$) given the tail entity and the relation is more straightforward compared to the reverse. We infer that the asymmetry stems from the more distinguishable spatial distribution of POI and land types, making them easier to associate with specific road segments compared to predicting a road based on its surrounding urban functions. (2) Spatial link influences. A decline in MR with increasing link numbers suggests that additional spatial links bolster the $\mathbb{K}_{S}$’s ability to delineate interactions and relationships between roads, POIs, and land uses. To further explore the link influence, we also investigate the embedding performance for Link[9] and Link[12] by adding more spatial links. However, the inclusion of Link[9] and Link[12] does not yield a large performance enhancement beyond Link[6]. For TransE, the performance on the right-side prediction even deteriorates for Link[9] and Link[12]. Hence, Link[6] is recommended for an optimal parameter of $\mathbb{K}_{S}$, balancing performance gains and computational resource usage. (3) Spatial buffer selection. We find that Buffer[10-100] achieves the best MR scores across three-side predictions using both models, corroborating findings from Section IV-B. This particular buffer range encapsulates essential spatial relationships in $\mathbb{K}_{S}$ among entities, striking a balance where the context is rich without being overly broad.
@@
Similarly, we analyze the three-side MR values for $\mathbb{K}_{T}$ using KG2E and TransE methods in Figure 4(c, d). (1) Head and tail predictions. We also observed the imbalanced phenomenon of left-side and right-side link predictions in embedding $\mathbb{K}_{T}$. For KG2E and TransE with different temporal links, predicting the head generally yields worse MR scores than predicting the tail. However, this imbalance becomes less evident compared to $\mathbb{K}_{S}$, which suggests a more uniform structure of the temporal unit. (2) Temporal context dependency. The MR values demonstrate a consistent decrease as the time buffers increase from Past[10] to Past[60] for all cases of embedding $\mathbb{K}_{T}$. This pattern underscores the importance of including historical information to enhance link prediction accuracy in $\mathbb{K}_{T}$. Moreover, integrating more temporal links, i.e., Link[HDW], consistently bolsters MR scores for both models across all configurations, reinforcing the findings in Section IV-C.
@@ V-B Forecasting traffic speed in limited datasets
To explore CKG-GNN’s efficacy in scenarios with potentially limited data, we apply the KGGCN models to two small datasets in Table VI, i.e., one from a typical workday $V_{work}$ and the other from a rest day $V_{rest}$. The divergence in day types is expected to manifest itself in the model’s performance due to the inherent differences in traffic patterns between workdays and rest days. The absence of weekly and daily temporal links in these single-day datasets leads to that only hourly links in $\mathbb{K}_{T}$ are utilized.
In Table VI, the CKG-enhanced models display superiority in speed prediction for both workday and rest day according to the consistently lower average MAE and MAPE for 10-120 min compared to the baseline DCRNN model. This phenomenon showcases CKG’s ability to model domain-specific context information for limited datasets, resulting in enhanced prediction accuracy. We find that a divergence in prediction ease between the workday and rest day is apparent. The $\mathbb{K}_{ST}$ model in 10-120 min demonstrates a lower MAE for the workday at 4.07±0.04 compared to 4.38±0.17 for the rest day. This discrepancy is attributed to the more structured traffic patterns typically observed during workdays, driven by routine commutes and schedules. Interestingly, when evaluating the incremental benefits offered by CKG-enhanced models, the results indicate a more pronounced improvement for the rest day than the workday. In detail, the average MAE improvement for $V_{work}$ and $V_{rest}$ is 0.16 (4.23-4.07) and 0.41 (4.79-4.38), respectively. This suggests that the integration of CKG is particularly effective in improving the less predictable and more variable traffic patterns that occur on rest days. On the rest day, the absence of regular commuting behaviors introduces complexities that context-aware knowledge graphs can offset, leading to a higher margin of improvement over the baseline.
TABLE VI: Prediction evaluation of CKG-enhanced models for two limited speed datasets. (w) refers to the workday dataset for $V_{work}$, and (r) refers to the rest day dataset for $V_{rest}$. Each experiment was run five times to obtain its mean and standard deviation.
Model 10 min 60 min 120 min Ave (10-120 min) MAE MAPE MAE MAPE MAE MAPE MAE MAPE (w)DCRNN 3.10±0.04 13.78±0.13 4.11±0.03 17.12±0.05 5.28±0.06 19.59±0.23 4.23±0.04 17.32±0.09 (w)CKG-$\mathbb{K}_{S}$ 3.08±0.01 13.51±0.05 4.10±0.01 17.00±0.03 5.13±0.04 19.41±0.14 4.18±0.02 17.15±0.05 (w)CKG-$\mathbb{K}_{T}$ 3.05±0.01 13.47±0.10 4.05±0.05 16.98±0.06 5.23±0.16 19.75±0.43 4.16±0.06 17.20±0.12 (w)CKG-$\mathbb{K}_{ST}$ 3.14±0.03 14.15±0.32 3.99±0.04 17.36±0.44 4.90±0.11 19.12±0.23 4.07±0.04 17.40±0.36 (r)DCRNN 3.27±0.06 14.01±0.51 4.51±0.04 16.33±0.05 6.46±0.14 20.33±0.29 4.79±0.05 16.97±0.12 (r)CKG-$\mathbb{K}_{S}$ 3.19±0.02 13.06±0.04 4.52±0.03 16.29±0.07 6.11±0.05 19.49±0.13 4.70±0.03 16.62±0.08 (r)CKG-$\mathbb{K}_{T}$ 3.14±0.05 12.73±0.06 4.35±0.09 15.97±0.18 5.89±0.15 19.15±0.45 4.54±0.09 16.27±0.21 (r)CKG-$\mathbb{K}_{ST}$ 3.16±0.04 13.16±0.28 4.27±0.18 15.88±0.24 5.43±0.28 17.88±0.55 4.38±0.17 16.00±0.24
Despite showing promise in limited datasets, the CKG-enhanced models also face some challenges when comparing performance across different forecast horizons. First, CKG-$\mathbb{K}_{S}$ and CKG-$\mathbb{K}_{T}$ individually outperform the baseline DCRNN model for 10-min prediction in the workday, but CKG-$\mathbb{K}_{ST}$ does not maintain this trend. We assume that when context information is focused—either spatially with $\mathbb{K}_{S}$ or temporally with $\mathbb{K}_{T}$—the models are able to effectively leverage this information for short-term predictions. When both types of context information are combined in $\mathbb{K}_{ST}$, the increased complexity requires more data to effectively train and refine its predictive capabilities in the brief 10-min forecasting interval. Second, $\mathbb{K}_{ST}$’s average performance in 10-120 min shows an improved MAE over the baseline for $V_{work}$, but its MAPE is not as favorable. The possible reasons are two-fold, i.e., relative error amplification during peak hours and lack of cyclical linkages in the single-day dataset. Since MAPE calculates the percentage error relative to the actual values, it amplifies errors when the base speed values are low in the workday peak hours. Also, MAPE’s sensitivity to errors emphasizes the challenge of accurately predicting traffic speed during congested periods, especially when lacking sufficient data samples to capture the cyclic nature of workday traffic patterns.

### v2

@@ IV-E Dual-view feature importance
Figure 4(a) demonstrates the heatmap derived from context-view attention weights in Equation 12 to investigate the contribution of context features. For the spatial unit, the prominence of the relation-dependent road feature at $[1]_{a}$ stands out, surpassing the influence of the road-only spatial feature at $[0]_{a}$ and the road-only temporal feature at $[10]_{a}$. This phenomenon suggests that isolated road representations without additional contextual information are less predictive. In examining the effectiveness of different spatial contexts, land uses at $[3]_{a}$ receive relatively higher importance scores compared to POIs at $[2]_{a}$. Despite this, both display less influence on traffic prediction compared to the relation-dependent road features at $[1]_{a}$. This lesser influence can be attributed to their indirect association with traffic systems, in contrast to the direct connections exhibited by road networks. The relatively high importance between land uses at $[3]_{a}$ and jam factors at $[12]_{a}$ underscores the impact of land use on traffic congestion factors, thereby contributing to traffic predictions. Moreover, the weights in Rows $[4-8]_{a}$ are notably elevated, indicating that spatial links—ranging from proximate to extended connections—play a crucial role in capturing the spatial dependency affecting traffic prediction. These links also exhibit a strong interplay with temporal features, particularly time indicators at Col $[11]_{a}$ and weather conditions at Col $[13]_{a}$.
@@
Regarding the temporal unit, all features within $[11-13]_{a}$ manifest as critical. First, time indicators and weather conditions in $[11,13]_{a}$ demonstrate substantial weights individually and mutually, indicating their strong influence on traffic speed predictions. Their interaction with the relation-dependent road feature at Col $[1]_{a}$ is also of great importance. Second, the jam factor at $[12]_{a}$ stands out for its significant self-weight, aligning with the intuitive understanding that the congestion level is directly predictive of traffic speed with an immediate impact. This emphasizes the critical role of jam factors as a key context feature in traffic forecasting. Furthermore, hourly and weekly patterns in Rows $[14,16]_{a}$ carry high weights, whereas daily patterns in Row $[15]_{a}$ appear less influential. The prominence of hourly patterns can be attributed to their immediate relevance to the prediction timeframe. Meanwhile, weekly patterns demonstrate consistent regularity, capturing the cyclical nature of traffic flow where behaviors on a specific day, like Monday, tend to mirror those from the previous week. In contrast, daily patterns exhibit a weaker correlation, suggesting that the connection between consecutive days, such as Friday to the following Monday, possesses limited predictive value in this context.
@@ V-A CKG embedding analysis
Regarding $\mathbb{K}_{S}$, we have several findings. (1) Head and tail predictions. An imbalance in MR values for left-side and right-side evaluations is observed for both methods in Figure 5(a, b), with the both-side case presenting intermediate values. This imbalance reveals that predicting the head entity (e.g., $poiType$ and $landType$) given the tail entity and the relation is more straightforward compared to the reverse. We infer that the asymmetry stems from the more distinguishable spatial distribution of POIs and land types, making them easier to associate with specific road segments compared to predicting a road based on its surrounding urban functions. (2) Spatial link influences. A decline in MR with increasing link numbers suggests that additional spatial links bolster the $\mathbb{K}_{S}$’s ability to delineate interactions and relationships between roads, POIs, and land uses. This trend is consistently observed across all buffer distances, indicating that denser connections within the transportation network can enhance the embedding performance of $\mathbb{K}_{S}$. To further explore the link influence, we also investigate the embedding performance for Link[9] and Link[12] by adding more spatial links. However, the inclusion of Link[9] and Link[12] does not yield a large performance enhancement beyond Link[6]. For TransE, the performance on the right-side prediction even deteriorates for Link[9] and Link[12]. Hence, Link[6] is recommended for an optimal parameter of $\mathbb{K}_{S}$, balancing performance gains and computational resource usage. (3) Spatial buffer selection. We find that Buffer[10-100] achieves the best MR scores across three-side predictions using both models, corroborating findings from Section IV-B. This particular buffer range encapsulates essential spatial relationships in $\mathbb{K}_{S}$ among entities, striking a balance where the context is rich without being overly broad.
@@
Similarly, we analyze the three-side MR values for $\mathbb{K}_{T}$ using KG2E and TransE methods in Figure 5(c, d). (1) Head and tail predictions. We also observed the imbalanced phenomenon of left-side and right-side link predictions in embedding $\mathbb{K}_{T}$. For KG2E and TransE with different temporal links, predicting the head generally yields worse MR scores than predicting the tail. However, this imbalance becomes less evident compared to $\mathbb{K}_{S}$, which suggests a more uniform structure of the temporal unit. (2) Temporal context dependency. The MR values demonstrate a consistent decrease as the time buffers increase from Past[10] to Past[60] for all cases of embedding $\mathbb{K}_{T}$. This pattern underscores the importance of including historical information to enhance link prediction accuracy in $\mathbb{K}_{T}$. Moreover, integrating more temporal links, i.e., Link[HDW], consistently bolsters MR scores for both models across all configurations, reinforcing the findings in Section IV-C. Both broadening the time buffer range from Past[10] to Past[60] and adding more temporal links from Link[-] to Link[HDW] introduce stronger temporal dependencies within $\mathbb{K}_{T}$, thereby improving the embedding effectiveness across various KGE techniques.
@@ V-B Forecasting traffic speed in limited datasets
To explore CKG-GNN’s efficacy in scenarios with potentially limited data, we apply the CKG-GNN models to two small datasets and investigate their traffic prediction performance in Table VI and Figure 6, i.e., one from a typical workday $V_{work}$ and the other from a rest day $V_{rest}$. We introduced the mean scaled interval score (MSIS) metric to quantify the uncertainty surrounding the point forecasts within a 95% prediction interval, as shown in Table VI. MSIS rewards models for providing narrow prediction intervals but penalizes them when actual values fall outside these intervals [33]. A lower MSIS indicates a more reliable uncertainty estimate, which is crucial for providing enhanced insights into the model’s confidence in its predicted outcome [34]. For details on MSIS computation, refer to Appendix-F. The divergence in day types is expected to manifest itself in the model’s performance due to the inherent differences in traffic patterns between workdays and rest days. The absence of weekly and daily temporal links in these single-day datasets leads to that only hourly links in $\mathbb{K}_{T}$ are utilized.
In Table VI, the CKG-enhanced models display superiority in 10-120 minute speed prediction for both workday and rest day according to the consistently lower average MAE, MAPE, and RMSE values compared to the baseline DCRNN model. This demonstrates that CKG can enhance the capacity of DCRNN to improve traffic forecasting accuracy, even with limited training data, through the integration of contextual information. The MSIS values also support this observation. CKG-$\mathbb{K}_{ST}$ achieved MSIS scores of 2.10 and 1.97, compared to 2.62 and 2.36 for DCRNN on the $V_{work}$ and $V_{rest}$ datasets, respectively, indicating lower uncertainty in the prediction intervals when integrating $\mathbb{K}_{ST}$. Since $V_{work}$ and $V_{rest}$ represent distinct traffic patterns, the improved accuracy, reflected in lower MAE, MAPE, RMSE, and the reduced uncertainty, indicated by lower MSIS, demonstrate that incorporating context-aware representations enables the model to learn domain-specific information, resulting in more accurate traffic forecasts despite the differences in traffic patterns.
In addition, when comparing the two limited datasets using $\mathbb{K}_{S}$, $\mathbb{K}_{T}$, and their combination $\mathbb{K}_{ST}$, we find that CKG-$\mathbb{K}_{ST}$ generally performs best, followed by CKG-$\mathbb{K}_{T}$, and then CKG-$\mathbb{K}_{S}$ for MAE, RMSE, and MSIS. The only exception is the MAPE for $V_{rest}$, where CKG-$\mathbb{K}_{S}$ yields the lowest value. This further verifies the observation in Table V that integrating temporal contexts into DCRNN typically leads to greater improvements than incorporating spatial contexts, even with limited datasets. Moreover, these results highlight the complementary strengths of spatial and temporal contexts. Spatial contexts enhance prediction accuracy by associating traffic systems with their surrounding environments, indirectly capturing human mobility patterns [8]. Conversely, temporal contexts directly capture the dynamic variations in traffic flow, making their influences more immediate [2]. The integration of both contexts allows for simultaneous leveraging of their characteristics, as shown in Figure 4, yielding superior results and validating the effectiveness of combining $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$ into $\mathbb{K}_{ST}$.
TABLE VI: Performance evaluation of CKG-enhanced models and DCRNN for traffic speed prediction (10–120 minutes) on two limited datasets.
Dataset Model MAE MAPE RMSE MSIS
$V_{work}$ DCRNN 4.79±0.05 16.97±0.12 6.52±0.06 2.62
CKG-$\mathbb{K}_{S}$ 4.70±0.03 16.62±0.08 6.48±0.05 2.49
CKG-$\mathbb{K}_{T}$ 4.54±0.09 16.27±0.21 6.18±0.10 2.46
CKG-$\mathbb{K}_{ST}$ 4.38±0.17 16.00±0.24 5.97±0.20 2.10
$V_{rest}$ DCRNN 4.23±0.04 17.32±0.09 5.94±0.05 2.36
CKG-$\mathbb{K}_{S}$ 4.18±0.02 17.15±0.05 5.93±0.02 2.25
CKG-$\mathbb{K}_{T}$ 4.16±0.06 17.20±0.12 5.83±0.08 2.28
CKG-$\mathbb{K}_{ST}$ 4.07±0.04 17.40±0.36 5.65±0.04 1.97
We then present the MAE, MAPE, and RMSE values for traffic prediction over the 10-120 minute range for a more detailed comparison in Figure 6. The central line represents the average performance, while the surrounding margin illustrates its standard deviation. The performance metrics across different time horizons align with the findings in Table VI, demonstrating that CKG-$\mathbb{K}_{ST}$ outperforms the other models in most cases. For $V_{work}$, the CKG-$\mathbb{K}_{ST}$ model consistently achieves the lowest MAE, MAPE, and RMSE values across the majority of time horizons in Figure 6(a1, a2, a3), while DCRNN produces higher errors throughout the 10-120 minute range. Notably, the reduction in MAE, MAPE, and RMSE between CKG-$\mathbb{K}_{ST}$ and DCRNN becomes more pronounced as the prediction horizon extends from 10 to 120 minutes, suggesting that the integration of CKG becomes especially effective for longer-term predictions. Similarly, for $V_{rest}$, CKG-$\mathbb{K}_{ST}$ remains the best-performing model in terms of MAE and RMSE in Figure 6(b1, b3), apart from MAPE in Figure 6(b2). The trend of greater error reduction as the time horizons extend from 10 to 120 minutes is also observed in this dataset.
As previously noted, MAPE is an exception for $V_{rest}$, where the integration of $\mathbb{K}_{ST}$ proves less effective compared to its performance in MAE and RMSE. Specifically, CKG-$\mathbb{K}_{ST}$ exhibits higher MAPE values during the 10-70 minute horizon but achieves lower values from 90 to 120 minutes. In contrast, CKG-$\mathbb{K}_{ST}$ outperforms other models in terms of MAE and RMSE across most time horizons. This discrepancy stems from the distinct ways MAE, MAPE, and RMSE handle errors: MAE treats all errors equally to measure overall bias, MAPE is highly sensitive to small actual speed values, and RMSE places more emphasis on larger errors. Thus, the strong performance in MAE and RMSE highlights CKG-$\mathbb{K}_{ST}$’s ability to reduce overall and large errors. However, MAPE underperforms due to its sensitivity to low actual speed values, where small base speeds can result in disproportionately high percentage errors. This makes MAPE less reliable in this scenario. Moreover, the lack of cyclical patterns in the single-day dataset, such as daily and weekly trends, also complicates traffic speed prediction.
Figure 6: Performance evaluation (ave±std) of CKG-enhanced models and DCRNN for traffic speed prediction from 10 to 120 minutes. (a1-a3) MAE, MAPE, and RMSE for $V_{work}$. (b1-b3) MAE, MAPE, and RMSE for $V_{rest}$.
V-C Computational cost and practical implications
Integrating context-aware knowledge graphs into the GNN model introduces additional parameters, thereby increasing the model’s complexity. To illustrate this, we present the parameter numbers and computational costs for our CKG-GNN model compared to the baseline DCRNN model in Table VII. The CKG-GNN model includes 2291.7K parameters, significantly more than the 372.4K parameters of the DCRNN model. However, the improvements in prediction accuracy justify these changes. For example, using the $V_{raw}$ dataset, the CKG-GNN model achieves reductions in MAE for 10-minute and 120-minute predictions by 0.06 and 0.26, respectively. Moreover, with limited datasets of $V_{work}$ and $V_{rest}$, CKG-GNN significantly enhances MAE by 1.03 in Figure 6(a1) and 0.38 in Figure 6(b1) for 120-minute predictions compared to DCRNN. Regarding computational costs, benchmarking experiments were conducted on a machine equipped with a single RTX 4090 GPU and a 16-core CPU. We distinguish two main costs here. First, embeddings of $\mathbb{K}_{S}$ and $\mathbb{K}_{T}$ required 911.1s and 105.6s, respectively, performed only once without further computational overhead. Second, training times for CKG-GNN are longer than for DCRNN due to higher parameter counts, specifically 19485.8s compared to 3746.2s. This trade-off is mitigated by the minimal impact on inference times, with CKG-GNN requiring 7.5s versus 1.7s for DCRNN. Despite the slightly longer inference time for CKG-GNN, the duration is manageable and practical for real-world applications.
TABLE VII: Parameter numbers and computation costs for DCRNN and CKG-GNN using $V_{raw}$. ’-’ indicates not applicable.
Method
#Params
Embed
time (s)
Train
time (s)
Inference
time (s)
DCRNN 372.4 - 3746.2 1.7
CKG Embed $\mathbb{K}_{S}$ 108.5 911.1 - -
Embed $\mathbb{K}_{T}$ 99.1 105.6 - -
CKG-GNN 2291.7 - 19485.8 7.5
The proposed method, despite its increased complexity, demonstrates strong generality through its integration with various publicly accessible context datasets, such as the free traffic API provided by HERE Technologies and the OSM platform. This accessibility facilitates the application of our framework to other urban settings. Furthermore, the framework is designed to incorporate additional types of contextual information, which could potentially enhance the CKG framework’s predictive capabilities by including broader contextual variables. In this case, although new entities and relations would be integrated into the CKG, the core elements of our framework would continue to function effectively. Moreover, the proposed CKG framework offers scalability through increased data volumes without affecting the CKG structure. Temporal data expansions, such as extending time frames, do not alter the CKG’s temporal unit structure, avoiding the need for re-embedding entities and relations. Spatial expansions, which might involve modeling multiple adjacent cities, are managed by segmenting study areas according to natural city boundaries defined by geographical and infrastructural characteristics. This potential for further scalability highlights a promising direction for future research.
In real-world scenarios, the integration of the CKG-GNN model into traffic management systems markedly improves our ability to forecast traffic speeds with high accuracy and efficiency. The CKG-GNN model not only surpasses baseline GNN models in predictive accuracy but also maintains a practical inference time of approximately 7 seconds during congested periods, such as the morning peak in $V_{raw}$. The increased accuracy and manageable inference time underscore the model’s suitability for real-time applications within traffic management systems, facilitating enhanced operational capabilities at traffic control centers. Specifically, it allows these centers to dynamically adjust traffic signals, optimize routes, and effectively manage traffic flows. Moreover, the CKG framework deepens our understanding of the dynamic interplay between various contextual factors and traffic systems, providing essential insights for refining traffic management strategies. By including diverse spatial contexts, such as POIs and land uses, the framework empowers traffic management systems to accurately assess traffic conditions under varying land use scenarios. The CKG’s temporal dimension illuminates the impacts of traffic congestion and weather conditions on traffic systems, enabling traffic control centers to respond adaptively to sudden traffic jams or severe weather events. This functionality ensures that traffic management systems can leverage real-time, data-driven decisions to significantly enhance traffic regulation and control, optimizing overall traffic flow and safety.